write the decimal and percent form of the fraction 4/5 (its a fraction)

Found 2 solutions by user_dude2008, Mathtut:
Answer by user_dude2008(1862)

4/5 = 0.8 <--- decimal

4/5 = 0.8 => 0.8*100 = 80%

4/5 = 80% <--- percent

Answer by Mathtut(3670)

4/5=.8 in decimal form
:
percentage form is .8(100)=80%